OpenStudy (gokuporter):

4(x-3)+5x=-48 A.all real numbers B.-4 C.0 d.4

OpenStudy (gokuporter):

A

OpenStudy (alexandervonhumboldt2):

first distribute: 4*(x-3)+5x=-48 4x-12+5x=-48 9x=-48+12 9x=-36 x=-4

OpenStudy (alexandervonhumboldt2):

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Gokuporter A \(\color{#0cbb34}{\text{End of Quote}}\) A is not correct

OpenStudy (alexandervonhumboldt2):

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Gokuporter A \(\color{#0cbb34}{\text{End of Quote}}\) A is not correct